OpenStudy (anonymous):

Relative maximum? g(x)=7/(x^2+5) . last question. Thanks for any help!

jimthompson5910 (jim_thompson5910):

assuming you know how to work with derivatives, you would derive it to get g(x)=7/(x^2+5) g(x)=7*(x^2+5)^(-1) g ' (x)=7*(-1)*(x^2+5)^(-2) * d/dx[x^2+5) g ' (x)=-7/( (x^2+5)^2 ) * (2x) g ' (x)=-14x/( (x^2+5)^2 )

jimthompson5910 (jim_thompson5910):

then you'd set the derivative equal to zero and solve for x once you figure out that x value, plug that x value back into g(x) to find the max value

OpenStudy (anonymous):

So should I set it up as -14x/((x^2+5)^2)=0 ? @jim_thompson5910